OpenStudy (anonymous):

Find the approximate value of the circumference of a circle with the given radius. Use = 3.14. Round your results to one more decimal than in the given radius. 4.6 inches C =

OpenStudy (alexandervonhumboldt2):

the formula for the curcimfence is 2*pi*r=2*3.14*4.6=solve

OpenStudy (anonymous):

thnx @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

np